# Flow Straighter and Faster: Efficient One-Step Generative Modeling via MeanFlow on Rectified Trajectories

### v1

@@ Abstract
Flow-based generative models have recently demonstrated strong performance, yet sampling typically relies on expensive numerical integration of ordinary differential equations (ODEs). Rectified Flow enables one-step sampling by learning nearly straight probability paths, but achieving such straightness requires multiple computationally intensive reflow iterations. MeanFlow achieves one-step generation by directly modeling the average velocity over time; however, when trained on highly curved flows, it suffers from slow convergence and noisy supervision. To address these limitations, we propose Rectified MeanFlow, a framework that models the mean velocity field along the rectified trajectory using only a single reflow step. This eliminates the need for perfectly straightened trajectories while enabling efficient training. Furthermore, we introduce a simple yet effective truncation heuristic that aims to reduce residual curvature and further improve performance. Extensive experiments on ImageNet at $64^{2}$, $256^{2}$, and $512^{2}$ resolutions show that Re-MeanFlow consistently outperforms prior one-step flow distillation and Rectified Flow methods in both sample quality and training efficiency. Code is available at Code is available at https://github.com/Xinxi-Zhang/Re-MeanFlow.
Figure 1: Re-Meanflow (Ours) leverages the synergy of trajectory rectification and mean-velocity modeling, achieving the best compute–quality trade-off—reaching strong FID. This synergy yields efficiency and quality that neither rectification nor mean-velocity modeling can achieve alone. All methods are initialized with pretrained EDM2-S [23].
@@ 1 Introduction
Flow models [27, 28] and diffusion models [43, 40] have become a central paradigm in generative modeling, enabling a wide range of applications across various data domains [18, 36, 55, 13, 14, 48]. Compared with earlier paradigms such as GANs [12, 20] and Normalizing Flows [35, 52], these models offer stable training and superior fidelity, but at the cost of expensive sampling. The root cause of this inefficiency is the curvature of the generative trajectories induced by the prior and data distributions. In practice, the velocity fields governing these flows bend sharply, making them difficult to approximate few discretization steps.
Rectified Flow [28, 24] addresses the multi-step sampling cost of flow models by progressively straightening trajectories through an iterative reflow process. In principle, if the trajectory becomes perfectly straight, one-step sampling is achievable. However, Liu et al. [28] theoretically show that achieving a fully straight trajectory requires infinitely many reflow iterations. More recently, Lee et al. [24] argue that a single reflow step is sufficient in practice, under the assumption that the intermediate flow mapping is approximately $L$-Lipschitz and that edge cases leading to curved trajectories are exceedingly rare. Although this assumption may hold for most data-noise couplings, we find that in practice noticeable edge cases still exist; for these cases, performing Euler one-step sampling on such curved paths will generate invalid samples.
In contrast, MeanFlow [11] completely bypasses ODE integration by directly modeling the mean velocity field over time, eliminating the need for perfectly straight paths and achieving strong single-step performance. However, for highly curved trajectories, MeanFlow suffers from unstable and noisy supervision signals, making the training slow and computationally expensive.
To address these limitations, we propose Re-Meanflow, short for Rectified MeanFlow, a conceptually simple yet training-effective framework that integrates the MeanFlow objective into the Rectified-flow framework. Specifically, we learn the mean velocity field along the rectified trajectory obtained after a single reflow step. This bypasses the need for perfectly straight trajectories while making MeanFlow training significantly more efficient and stable, as the rectified trajectory exhibits substantially lower variance than the original highly curved flow. Additionally, we introduce a simple and surprisingly effective truncation heuristic that reduces the influence of extreme curvature cases during training. Instead of attempting to explicitly detect curvature, we discard the top $10\%$ couplings ranked by $\ell_{2}$ trajectory distance. While this heuristic does not guarantee that all high-curvature trajectories are removed, we hypothesize that large distances often correlate with more curved paths, and applying this filter consistently improves both training stability and sample quality in practice.
@@
Extensive experiments on ImageNet at $64^{2}$, $256^{2}$, and $512^{2}$, demonstrating that Re-Meanflow consistently outperforms both state-of-the-art distillation methods and strong train-from-scratch baselines in terms of both generation quality and training efficiency. In particular, compared to its closely related counterpart, 2-rectified flow++ [24], Re-Meanflow reduces FID by 33.4% while requiring only 10% of the total compute. Beyond empirical gains, Re-Meanflow introduces a practical shift in how few-step generative models can be trained. Existing distillation pipelines depend heavily on high-end training GPUs, making hyperparameter tuning and repeated runs prohibitively expensive. In contrast, Re-Meanflow offloads the majority of computation to the inference-driven reflow stage, which can be executed on widely available consumer- or inference-grade accelerators, and requires only a lightweight MeanFlow training phase. The training stage of Re-Meanflow accounts for merely 17% of the total GPU hours used by AYF[37].
2 Related Works
@@
Diffusion and Flow Matching. Diffusion/flow-based generative models [40, 43, 17, 44, 27, 1, 28] learn to reverse a gradual noising process, where the reverse-time dynamics can be formulated as a deterministic probability-flow ODE [44, 21]. Although these approaches achieve strong performance, they typically require multi-step numerical integration for sampling [41, 21, 30, 53] due to the high curvature of generative paths.
@@
Few-step Diffusion/Flow Models. Rectified Flow methods [28, 46, 24] alleviate this by explicitly learning straighter trajectories that enable one-step or few-step sampling, but such straight-path construction itself incurs high computational cost due to heavy training and repeated reflow procedures. Consistency Models [45, 42, 10, 29, 49] train networks to produce invariant outputs across different timesteps, enabling direct few-step or even one-step sampling. Flow Map Models [5, 9, 11] bypass ODE integration by learning the displacement or velocity map over time. Despite their strong empirical results, these few-step paradigms often face stability challenges or high training cost because they must learn mappings along inherently curved trajectories, where supervision is noisy, and optimization is difficult. Recent work has sought to mitigate this issue by simplifying the consistency objective [29] or introducing improved loss functions and normalization strategies[6]. Concurrent work CMT [19] aims to improve the efficiency of few-step models by supervising on learned ODE trajectories. Building on these insights, this work aims to combine trajectory simplification with one-step modeling.
@@
Efficient Training for Diffusion/Flow Models. Many recent works [51, 26, 50, 47] boost the training efficiency of diffusion models by leveraging external representation learners like DINO [33]. Orthogonal works [56, 38] adopt masked-transformer designs that exploit spatial redundancy by training on only a subset of tokens. ECT [10] enables efficient consistency models through a progressive training strategy that transitions from diffusion to consistency training. Re-Meanflow contributes in this direction by training a MeanFlow model on a significantly simplified flow path.
Figure 2: Why trajectory rectification and mean-velocity modeling reinforce each other. (a) A 1-rectified flow (Refer to caption) still follows highly curved trajectories, requiring many ODE steps. Applying two rounds of rectification (Refer to caption) straightens the paths and reduces NFEs, but one-step sampling (Refer to caption) remains unreliable unless trajectories are nearly straight. (b) MeanFlow estimates the average velocity $u(\mathbf{z}_{t},r,t)$ over all intervals $(r,t)$, which in principle avoids the need for straight paths. However, when the underlying velocity field is highly curved, the induced averages become complex and hard to learn. (c) Training MeanFlow on trajectories from a 2-rectified flow yields a significantly smoother average-velocity field, making estimation easier and enabling faster convergence and high-quality one-step generation.
3 Preliminaries
Rectified Flow[28]. Given a prior distribution $p_{\mathbf{z}}$ (taken as $\mathcal{N}(\mathbf{0},\mathbf{I})$ in this work) and a data distribution $p_{\mathbf{x}}$, Rectified Flow trains a flow model $v_{\theta}$ that transports $p_{\mathbf{z}}$ to $p_{\mathbf{x}}$ via an ordinary differential equation (ODE) over time $t\in[0,1]$. For a data-noise coupling $(\mathbf{x},\mathbf{z})\sim p_{\mathbf{x}\mathbf{z}}$, where $p_{\mathbf{x}\mathbf{z}}$ denotes the joint distribution of $\mathbf{x}$ and $\mathbf{z}$, the flow path is defined as a linear interpolation $\mathbf{z}_{t}=(1-t)\mathbf{x}+t\mathbf{z}$. The flow model $v_{\theta}$ is trained by regressing the conditional velocity of observed couplings:
@@ 3 Preliminaries
$\mathcal{L}_{\text{RE}}(\theta)=\mathbb{E}_{\mathbf{x},\mathbf{z}\sim p_{\mathbf{x}\mathbf{z}},\ t\sim p_{t}}\left[\left\lVert\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{z}_{t}-v_{\theta}(\mathbf{z}_{t},t)\right\rVert_{2}^{2}\right]\quad\text{with }\ \frac{\mathrm{d}}{\mathrm{d}t}\mathbf{z}_{t}=\mathbf{z}-\mathbf{x}$ (1)
@@
Once $v_{\theta}$ is learned, a new sample $\mathbf{x}$ can be generated by solving the ODE for $\mathbf{z}\sim p_{\mathbf{z}}$:
@@
Rectified Flow begins training with independently sampled $\mathbf{x}$ and $\mathbf{z}$, i.e., $p^{0}_{\mathbf{x}\mathbf{z}}(\mathbf{x},\mathbf{z})=p_{\mathbf{x}}(\mathbf{x})p_{\mathbf{z}}(\mathbf{z})$. Training on these independent couplings produces highly curved ODE trajectories, requiring a large number of function evaluations (NFEs) to accurately solve Eq. 2. To mitigate this, Rectified Flow adopts a recursive reflow process to iteratively straighten the ODE trajectories. Given a learned velocity field $v_{\theta}^{k}$ trained on $p^{k-1}_{\mathbf{x}\mathbf{z}}$, a straighter velocity field $v_{\theta}^{k+1}$ can be obtained by training on $p^{k}_{\mathbf{x}\mathbf{z}}$, which is sampled by the previous velocity field $v_{\theta}^{k}$. For clarity, we refer to the trained vector field $v_{\theta}^{k+1}$ as the (k+1)(k{+}1)-rectified flow.
MeanFlow [11]. Compared to Rectified Flow [28], which models the instantaneous velocity and requires solving the numerical ODE in Eq. 2 during sampling, MeanFlow enables one-step sampling by modeling the average velocity $u$ between two time points $t$ and $r$:
$u(\mathbf{z}_{t},r,t)\triangleq\frac{1}{t-r}\int_{r}^{t}v(\mathbf{z}_{\tau},\tau)d\tau.$ (3)
To train a a neural network $u_{\theta}(\mathbf{x}_{t},r,t)$ to approximate this average velocity, MeanFlow derives an implicit training target linking the average velocity, the instantaneous velocity $v(\mathbf{x}_{t},t)$, and the time derivative of $u_{\theta}$:
$\mathcal{L}_{MF}(\theta)=\mathbb{E}_{\mathbf{x},\mathbf{z},r,t}\left\lVert u_{\theta}(\mathbf{z}_{t},r,t)-\operatorname{sg}(u_{\text{tgt}})\right\rVert^{2}_{2}$ (4)
@@
Here, $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator, and the derivative $\tfrac{d}{dt}u(\mathbf{z}_{t},r,t)$ can be computed using the Jacobian-vector product (JVP) interface in frameworks such as PyTorch or JAX. once $u_{\theta}$ is trained, the entire flow path can be reconstructed from a single evaluation $u_{\theta}(\mathbf{z},0,1)$ without numerical integration.
4 Methods
To motivate Re-Meanflow, we first examine the limitations of existing rectified-flow and MeanFlow approaches and illustrate why combining them yields a substantial synergy with a toy example.
Figure 3: A Class-imbalanced 2D Toy Example. We consider a controlled 2D setup where a flow model transports a balanced two-component Gaussian mixture on the left to an imbalanced mixture on the right (weights $0.4$ for the upper mode and $0.6$ for the lower mode). Panels (a-c): (a) Linear interpolation of independently sampled couplings $p_{\mathbf{x}}\times p_{\mathbf{z}}$, which serve as the training signal for the first velocity model. (b) The resulting 1-rectified flow learned from these independent couplings; the learned velocity field remains noticeably curved. (c) Using the velocity field from (b), we generate a new set of couplings and train a second velocity model on their linear interpolations, yielding the 2-rectified flow. Panel (d): Due to imperfect straightening, one-step Euler sampling on the 2-rectified flow still yields noticeable outliers. Panel (e): MeanFlow trained directly on independent couplings fails to converge within the training budget because high-variance conditional velocities destabilize learning. Panel (f): Re-Meanflow combines trajectory rectification with MeanFlow, eliminating most outliers and achieving more accurate one-step generation.
2D Toy Example Setting. Figure 3 visualizes a simple transport task from a balanced two-mode Gaussian mixture to an imbalanced two-mode mixture. We compare one-step generation performance across three approaches under a fixed training budget of 20k iterations. We evaluate:
  1. 1.
    2-rectified flow, where each reflow iteration is trained for 10k steps;
  2. 2.
    MeanFlow, trained for 20k steps directly on independently sampled couplings;
  3. 3.
    Re-Meanflow (ours), which first trains a velocity model for 10k steps to obtain a 1-rectified flow, followed by 10k steps of MeanFlow training on the resulting couplings.
More Detail of the toy example setting are provided in Appendix B.5.
4.1 Limitations of Rectified flow and MeanFlow
Rectified Flow Leaves Residual Curvature Degrading One-Step Sampling.
Rectified Flow seeks to enable one-step sampling by progressively straightening ODE trajectories. However, Liu et al. [28] show that achieving a perfectly straight trajectory theoretically requires applying an infinite number of reflow iterations. Later, Lee et al. [24] argue that a single reflow step already yields trajectories that are “straight enough” for one-step Euler sampling, since couplings in $p^{1}_{\mathbf{x}\mathbf{z}}$ that would induce curved paths are exceedingly rare under the assumption that the 1-rectified flow is $L$-Lipschitz:
$\|\mathbf{x}^{\prime}-\mathbf{x}^{\prime\prime}\|_{2}\leq L\|\mathbf{z}^{\prime}-\mathbf{z}^{\prime\prime}\|_{2}\quad\text{ where }(\mathbf{x}^{\prime},\mathbf{z}^{\prime}),(\mathbf{x}^{\prime\prime},\mathbf{z}^{\prime\prime})\sim p^{1}_{\mathbf{x}\mathbf{z}}.$ (6)
While the $L$-Lipschitz condition guarantees that nearby noise samples cannot map to arbitrarily distant data points, it does not prevent $L$ from being large in practice. When $L$ is large, even modest differences in noise space can translate into substantial separation in data space, producing early-time bending in the transport paths and degrading the accuracy of one-step Euler sampling.
Toy Evidence (Panel d). Due to the effect of class imbalance, samples associated with the lighter target mode must traverse disproportionately long distances toward the denser component. Consequently, even when their noise samples are close, the corresponding data points can lie far apart, yielding large effective $L$ in Eq. 6 and causing trajectories to bend near $t\!\approx\!0$. The severe one-step deviations in Fig. 3(d) reflect precisely this early-time curvature. Similar geometric imbalances arise naturally in real datasets, underscoring the need for methods that remain stable and accurate even in the presence of residual curvature after rectification.
MeanFlow Struggles on Highly Curved Independent Couplings. Training MeanFlow directly on curved flows suffers from noisy and unstable supervision signals, which is a well-known cause of slow convergence in machine learning. Concretely, in the training objective of Eq. 4, the marginal velocity field is not directly available and must be approximated from conditional velocities. Because the couplings are sampled independently, these conditional velocities exhibit high variance. Moreover, the derivative term $(t-r)\tfrac{d}{dt}u(\mathbf{z}_{t},r,t)$ relies on Jacobian-vector products of the model output without explicit supervision, introducing additional noise that compounds the variance. Finally, when trajectories are highly curved, the corresponding average velocity field becomes irregular and difficult to approximate, as illustrated in Fig. 2b.
Toy Evidence (Panel e). MeanFlow trained directly on independent couplings struggles to learn a coherent transport map under the same training budget, resulting in poor overall performance.
4.2 Re-Meanflow: Efficient One-Step Generative Modeling via MeanFlow on Rectified Trajectories

### v2

# Overcoming the Curvature Bottleneck in MeanFlow

@@ Abstract
MeanFlow offers a promising framework for one-step generative modeling by directly learning a mean-velocity field, bypassing expensive numerical integration. However, we find that the highly curved generative trajectories of existing models induce a noisy loss landscape, severely bottlenecking convergence and model quality. We leverage a fundamental geometric principle to overcome this: mean-velocity estimation is drastically simpler along straight paths. Building on this insight, we propose Rectified MeanFlow, a self-distillation approach that learns the mean-velocity field over a straightened velocity field, induced by rectified couplings from a pretrained model. To further promote linearity, we introduce a distance-based truncation heuristic that prunes residual high-curvature pairs. By smoothing the optimization landscape, our method achieves strong one-step generation performance. We improve the FID of baseline MeanFlow models from 30.9 to 8.6 under same training budget, and outperform the recent 2-rectified flow++ by 33.4% in FID while running 26x faster. Our work suggests that the difficulty of one-step flow generation stems partially from the rugged optimization landscapes induced by curved trajectories. Code is available at https://github.com/Xinxi-Zhang/Re-MeanFlow.
Figure 1: Straightening flows smooths the loss landscape for one-step generation. (a) MeanFlow exhibits a sharply peaked and irregular loss landscape, making optimization difficult for learning efficient one-step generators. (b) By learning the mean velocity on a rectified coupling with substantially straighter trajectories, Re-MeanFlow (Ours) yields a much smoother and more regular objective. (c) This improved landscape empirically leads to faster convergence and stronger one-step generation, even when MeanFlow is trained for $2\times$ longer. All plots are from ImageNet $256^{2}$, with both MeanFlow and Re-MeanFlow initialized from the same SiT-XL model.
@@ 1 Introduction
Flow models [lipman2022flow, liu2022flow] and diffusion models [song2019generative, sohl2015deep] have become a central paradigm in generative modeling, enabling a wide range of applications across various data domains [ho2022video, rombach2022high, zhang2025soda, he2023dmcvr, he2024dice, xu2025can]. Compared with earlier paradigms such as GANs [goodfellow2020generative, karras2019style] and Normalizing Flows [rezende2015variational, zhai2024normalizing], these models offer stable training and superior fidelity, but at the cost of expensive sampling. The root cause of this inefficiency is the curvature of the generative trajectories induced by the mismatch between the prior and data distributions. In practice, the resulting velocity field can bend sharply, making it difficult to approximate accurately with only a few discretization steps.
This drawback has motivated works [liu2022flow, tong2023improving, onken2021ot, liu2022rectified] to straighten the underlying trajectories via optimal transport; however, existing methods do not produce sufficiently linear paths, which in turn prevents reliable one-step sampling based on the instantaneous velocity. Another line of work [song2023consistency, kim2023consistency, frans2024one, geng2025mean, zhou2025inductive] bypasses ODE integration by directly learning time-indexed flow-map predictions. Among these approaches, MeanFlow [geng2025mean] is particularly appealing: it models the time-dependent mean-velocity field, enabling single-step generation without requiring straight trajectories, and achieves strong empirical performance. Despite this advantage, training MeanFlow remains costly and often converges slowly, even when initialized from a pretrained flow model.
In this work, we identify a key bottleneck in mean-velocity learning: the curvature of the underlying generative trajectories. Motivated by this insight, we introduce Rectified MeanFlow (Re-MeanFlow), a lightweight self-distillation approach that trains MeanFlow on rectified couplings, yielding markedly straighter trajectories and a simpler mean-velocity learning problem. Fig. 1 highlights this contrast: optimizing MeanFlow on the original (unrectified) trajectories produces a highly rugged landscape with sharp spikes, while Re-MeanFlow yields a noticeably smoother and better-conditioned surface. This improved conditioning leads to substantially faster convergence and superior one-step generation, even compared against MeanFlow trained with $2\times$ more compute. Importantly, Re-MeanFlow is data-free: it requires only a pretrained flow model and samples from the prior to generate rectified couplings, without access to the original training dataset.
Additionally, we employ a simple distance-based truncation heuristic to further reduce curvature during training. Motivated by the empirical correlation between trajectory curvature and endpoint distance, we discard the top $10\%$ of couplings ranked by their $\ell_{2}$ distance between endpoints. This filter removes residual high-curvature trajectories and consistently improves both training stability and sample quality in our experiments.
Extensive experiments on ImageNet at $64^{2}$, $256^{2}$, and $512^{2}$ show that Re-MeanFlow consistently outperforms state-of-the-art distillation methods and strong train-from-scratch baselines in both generation quality and training efficiency. In particular, relative to the closely related 2-rectified flow++ [lee2024improving], Re-MeanFlow reduces FID by 33.4% while using only 10% of the compute.
@@
Beyond empirical gains, Re-MeanFlow suggests a more practical training paradigm for few-step generative models. Existing distillation pipelines rely heavily on high-end training GPUs, making hyperparameter tuning and repeated runs prohibitively expensive. In contrast, Re-MeanFlow shifts most computation to an inference-driven reflow stage that can be executed on widely available consumer- or inference-grade accelerators, followed by a lightweight MeanFlow training phase. Overall, the training cost in Re-MeanFlow accounts for only 17% of the total GPU hours used by AYF [sabour2025align].
2 Related Work
We review diffusion and flow matching, emphasizing methods that reduce sampling cost via trajectory straightening or few-step modeling, and summarize recent advances in efficient training.
2.1 Diffusion and Flow Matching.
@@ 2 Related Work
Diffusion/flow-based generative models [sohl2015deep, song2019generative, ho2020denoising, song2020score, lipman2022flow, albergo2022building, liu2022flow] learn to reverse a gradual noising process, where the reverse-time dynamics can be formulated as a deterministic probability-flow ODE [song2020score, karras2022elucidating]. Although these approaches achieve strong performance, they typically require multi-step numerical integration for sampling [song2020denoising, karras2022elucidating, lu2022dpm, zhang2022fast] due to the high curvature of generative paths.
2.2 Straightening the Generative Trajectories.
The high curvature of diffusion/flow generative trajectories has motivated work that explicitly reduces the transport curvature between noise and data. OT-Flow [onken2021ot] regularizes continuous normalizing flows with OT-inspired objectives to encourage straighter dynamics. OT-CFM [tong2023improving] and Multisample Flow Matching [pooladian2023multisample] further reduce curvature by introducing OT-like minibatch couplings, yielding simpler probability paths and lower-variance supervision. However, such minibatch couplings are inherently local and do not directly enforce globally straighter trajectories. In contrast, Rectified Flow [liu2022flow] achieves global trajectory straightening via reflow, iteratively refining the transport to reduce transport cost and producing near-linear paths that admit accurate few-step simulation. We therefore adopt rectified trajectories as a clean, low-curvature target for MeanFlow training, substantially easing one-step optimization.
2.3 Few-step Diffusion/Flow Models.
@@
Few-step models aim to bypass numerical integration by directly learning large time-step transitions. Consistency Models [song2023consistency, song2023improved, geng2024consistency, lu2024simplifying, yang2024consistency] train networks to produce invariant outputs across different timesteps, enabling direct few-step or even one-step sampling. Flow Map Models [boffi2024flow, frans2024one, geng2025mean] bypass ODE integration by learning the displacement or velocity map over time. Despite their strong empirical results, these few-step paradigms often face stability challenges or high training cost because they must learn mappings along inherently curved trajectories, where supervision is noisy, and optimization is difficult. Recent work has sought to mitigate this issue by simplifying the consistency objective [lu2024simplifying] or introducing improved loss functions and normalization strategies[dao2025improved]. Concurrent work CMT [hu2025cmt] improves the efficiency of few-step models by supervising them with teacher ODE trajectories. Building on these insights, this work aims to combine trajectory simplification with one-step modeling.
2.4 Efficient Training for Diffusion/Flow Models.
@@
Many recent works [yu2024representation, leng2025repa, yao2025reconstruction, wu2025representation] boost the training efficiency of diffusion models by leveraging external representation learners such as DINO [oquab2023dinov2]. Orthogonal works [zheng2023fast, sehwag2025stretching] adopt masked-transformer designs that exploit spatial redundancy by training only on a subset of tokens. ECT [geng2024consistency] enables efficient consistency models through a progressive training strategy that transitions from diffusion to consistency training. Re-MeanFlow contributes in this direction by allowing efficient one-step modeling on a significantly simplified flow path.
3 Efficient Mean-Velocity Modeling on Straightened Trajectory
In this section, we first review one-step flow-based generation via MeanFlow (Sec. 3.1). We then identify the curvature bottleneck: the standard independent coupling induces highly curved paths that hinder mean-velocity learning (Sec. 3.2). Finally, we introduce Re-MeanFlow, which models mean-velocity on rectified, substantially straighter trajectories, and describe our practical training pipeline, including CFG and distance-based truncation (Sec. 3.4).
3.1 One-Step Flow-Based Generative Modeling via MeanFlow
Figure 2: Method. (a) MeanFlow learns the mean-velocity $u(\mathbf{z}_{t},r,t)$ on highly curved trajectories, which makes optimization difficult. In contrast, (b) Re-MeanFlow (ours) models $u(\mathbf{z}_{t},r,t)$ on a rectified coupling that induces substantially straighter generative trajectories, simplifying the underlying vector field and enabling faster and more stable training. $\mathrm{Curv}(r,t)$ denotes the curvature proxy measuring trajectory curvature between time points $r$ and $t$ (see Sec. 3.3).
Flow-based generative models transport a prior $p_{\mathbf{z}}$ (taken as $\mathcal{N}(\mathbf{0},\mathbf{I})$ in this work) to the data distribution $p_{\mathbf{x}}$ by learning a time-dependent velocity field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that induces a path of intermediate densities $\{p_{t}\}_{t\in[0,1]}$ in $\mathbb{R}^{d}$. Concretely, given a data-noise coupling $p_{\mathbf{x}\mathbf{z}}$, let $(\mathbf{x},\mathbf{z})\sim p_{\mathbf{x}\mathbf{z}}$ and define the linear interpolation as:
$\mathbf{z}_{t}=(1-t)\mathbf{x}+t\mathbf{z}.$
Let $p_{t}$ denote the distribution of $\mathbf{z}_{t}$. The coupling $p_{\mathbf{x}\mathbf{z}}$ together with the interpolation, therefore, specifies the probability path $\{p_{t}\}_{t\in[0,1]}$.
The associated velocity field $v(\mathbf{z}_{t},t)$ is defined as the vector field whose continuity equation
$\partial_{t}p_{t}+\nabla\cdot(p_{t}v(\cdot,t))=0$
matches this density evolution. Along the interpolation trajectory, the instantaneous velocity is given by
$\frac{d}{dt}\mathbf{z}_{t}=\mathbf{z}-\mathbf{x}.$
Flow Matching (FM) [lipman2022flow, liu2022flow] models the time-dependent velocity field with a neural vector field $v_{\theta}(\cdot,t)$ and trains it by regressing the trajectory velocity induced by the chosen coupling:
@@ 3.1 One-Step Flow-Based Generative Modeling via MeanFlow
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{(\mathbf{x},\mathbf{z})\sim p_{\mathbf{x}\mathbf{z}},\ t\sim p(t)}\left[\left\lVert\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{z}_{t}-v_{\theta}(\mathbf{z}_{t},t)\right\rVert_{2}^{2}\right]$ (1)
@@
Once $v_{\theta}$ is learned, a new sample $\mathbf{x}$ can be generated by solving the ODE for $\mathbf{z}\sim p_{\mathbf{z}}$ using a numerical solver:
@@
MeanFlow [geng2025mean] instead parameterizes the mean-velocity between two time points $r<t$:
$u(\mathbf{z}_{t},r,t)\triangleq\frac{1}{t-r}\int_{r}^{t}v(\mathbf{z}_{\tau},\tau)\,d\tau,$ (3)
To train a neural network $u_{\theta}(\mathbf{z}_{t},r,t)$ to approximate this mean-velocity, MeanFlow derives an implicit training target that connects the mean-velocity, the instantaneous velocity $v(\mathbf{z}_{t},t)$, and the time derivative of $u_{\theta}$:
$\mathcal{L}_{MF}(\theta)=\mathbb{E}_{(\mathbf{x},\mathbf{z})\sim p_{\mathbf{x}\mathbf{z}},\,(r,t)\sim p_{r,t}}\left\lVert u_{\theta}(\mathbf{z}_{t},r,t)-\operatorname{sg}(u_{\text{tgt}})\right\rVert^{2}_{2}$ (4)
@@
Here, $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator. Once $u_{\theta}$ is trained, sampling reduces to a single evaluation $u_{\theta}(\mathbf{z},0,1)$, avoiding the numerical integration required by Flow Matching (cf. Eq. 2).
3.2 Taming the Curvature Bottleneck
Crucially, the underlying generative trajectories of MeanFlow are dictated by the data-noise coupling distribution $p_{\mathbf{x}\mathbf{z}}$. In the absence of prior knowledge about how data and noise should be paired, MeanFlow (and many flow-based methods) adopts the independent coupling distribution $p^{0}_{\mathbf{x}\mathbf{z}}(\mathbf{x},\mathbf{z})=p_{\mathbf{x}}(\mathbf{x})\,p_{\mathbf{z}}(\mathbf{z}),$ which is known to induce highly curved generative trajectories (Fig. 2a).
We argue that learning the mean-velocity field on such curved trajectories is challenging: curvature amplifies the complexity of the target mean-velocity, which induces a rugged and poorly conditioned loss landscape (visualized in Fig. 1a and Fig. 7 (top)), amplifying the effect of imperfect supervision and slowing down optimization (see Sec. 4.3 for further analysis). To address this bottleneck, we propose Rectified MeanFlow (Re-MeanFlow), which models the mean-velocity field on a rectified coupling that induces substantially straighter trajectories (Fig. 2b). This rectification simplifies the underlying vector field and, correspondingly, yields a markedly smoother and more regular objective (Fig. 1b and Fig. 7 (bottom)), enabling faster convergence and improved one-step generation.